# Adaptive LoRA Experts Allocation and Selection for Federated Fine-Tuning

### v1

@@ C.1 Clustering Results
C.2 Sensitivity Analysis
@@
We vary the number of local training epochs to examine its effect on performance. The results in fig.˜9 show that our proposed method consistently outperforms all baseline methods across different local epoch settings ($E=5$), confirming that FedLEASE’s effectiveness is not dependent on specific epoch configurations.
Impact of LoRA Rank. We test the performance with different LoRA ranks, adjusting the rank parameter for both our method and baselines. As shown in table˜4, FedLEASE maintains superior performance across all tested rank values. The performance gap is particularly notable at lower ranks, highlighting our method’s efficiency in parameter utilization.
@@
Impact of Client Numbers. To evaluate scalability, we change the number of clients in the system to 8 and 32. table˜5 demonstrates that our proposed method maintains its performance advantage with different number of clients, indicating robust scalability to federated networks.
@@
Impact of Data Heterogeneity. We evaluate our method across varying degrees of data heterogeneity, focusing on realistic task heterogeneity rather than simple label distribution shifts. We consider three unbalanced task distribution settings: (1) Least heterogeneous: 16 clients having 2 kinds of NLU datasets (10 with QNLI and 6 with QQP); (2) Mildly heterogeneous: 16 clients having 3 kinds of NLU datasets (4 with SST-2, 7 with QNLI, 5 with QQP); and (3) Most heterogeneous: 16 clients having 4 kinds of NLU datasets (3 with SST2, 6 with QNLI, 2 with MRPC, 5 with QQP). The results in table˜6 demonstrate that FedLEASE consistently outperforms baseline methods across all heterogeneity levels, with the performance advantage becoming more pronounced as heterogeneity increases.
@@
We further conduct the additional experiments under label non-i.i.d. setting, and the label distribution is illustrated in fig.˜9. Note the task distribution is same as what we used in the main experiment. As shown in table˜7, our proposed mthod still outperforms other baselines in this label Non-IID setting.
Table 7: Performance Under Non-IID Label Distribution.
@@
Impact of Expert Upper Bound $M_{\max}$. In our main experiments, we set the expert upper bound to $M_{\max}=8$. To investigate the sensitivity of our approach to this parameter, we conducted additional experiments with lower upper bounds ($M_{\max}=2,3,4$) using the same experimental setup: 16 clients with data from 4 GLUE datasets. As shown in table˜8, we observe comparable performance between $M_{\max}=8$ and $M_{\max}=4$ configurations. This aligns with our clustering analysis in section˜C.1, which indicates that this particular configuration requires only 4 experts. This finding validates that our method can efficiently determine the appropriate number of experts needed even when the budget ($M_{\max}$) exceeds the system’s actual requirements. When restricted to $M_{\max}<4$, we observe a performance degradation compared to the $M_{\max}=4$ or $M_{\max}=8$ settings, further confirming the importance of allocating an adequate number of experts. Nevertheless, it is noteworthy that even with this constrained expert budget, our proposed method still outperforms all baseline methods. This demonstrates the robustness of our approach and its ability to make efficient use of even limited expert resources through effective allocation and adaptive selection.
Appendix D Limitations
@@
Appendix E Convergence Analysis
@@ E.1 Assumptions
Assumption E.1 (Client-Level Smoothness).
@@
Assumption E.2 (Client-Level Strong Convexity).
@@
Assumption E.3 (Bounded Expert Sensitivity).
@@
Assumption E.4 (Cluster Assignment Stability).
@@
Theorem E.5 (Convergence of FedLEASE).

### v2

@@ Appendix B Clustering Results
While clustering is indeed an essential component of our proposed method—playing a key role in expert allocation based on client similarity—the specific choice of clustering algorithm is not the focus of our contribution. Our observations in Section˜3.3 suggest that as long as the method captures pairwise similarity between clients (e.g., via LoRA B matrices), the overall performance is relatively robust to the particular clustering strategy.
We adopt Agglomerative Hierarchical Clustering due to its ability to operate directly on pairwise distances without requiring pre-defined centroids. To validate the generality of our approach, we also applied Spectral Clustering, which similarly supports pairwise similarity inputs, and observed comparable performance. Tables˜4 and 5 below demonstrate that both clustering methods achieve similar silhouette scores and downstream performance, reinforcing that our performance gains stem primarily from the expert allocation and adaptive top- selection mechanisms, rather than the specific clustering algorithm used.
Table 4: Silhouette Scores Comparison
Clustering Method 2 3 4 5 6 7 8
Spectral Clustering 0.0585 0.0820 0.1023 0.0739 0.0637 0.0549 0.0218
Agglomerative Hierarchical Clustering 0.0599 0.0884 0.1066 0.0965 0.0854 0.0747 0.0660
Table 5: Performance Comparison
Clustering Method SST2 QNLI MRPC QQP Average
Spectral Clustering 93.97 86.63 86.48 83.40 87.62
Agglomerative Hierarchical Clustering 93.33 87.22 86.93 83.57 87.76
Appendix C Sensitivity Analysis
@@
We vary the number of local training epochs to examine its effect on performance. The results in Figure˜9 show that our proposed method consistently outperforms all baseline methods across different local epoch settings ($E=5$), confirming that FedLEASE’s effectiveness is not dependent on specific epoch configurations.
Impact of LoRA Rank. We test the performance with different LoRA ranks, adjusting the rank parameter for both our method and baselines. As shown in Table˜6, FedLEASE maintains superior performance across all tested rank values. The performance gap is particularly notable at lower ranks, highlighting our method’s efficiency in parameter utilization.
@@
Impact of Client Numbers. To evaluate scalability, we change the number of clients in the system to 8 and 32. Table˜7 demonstrates that our proposed method maintains its performance advantage with different number of clients, indicating robust scalability to federated networks.
@@
Impact of Data Heterogeneity. We evaluate our method across varying degrees of data heterogeneity, focusing on realistic task heterogeneity rather than simple label distribution shifts. We consider three unbalanced task distribution settings: (1) Least heterogeneous: 16 clients having 2 kinds of NLU datasets (10 with QNLI and 6 with QQP); (2) Mildly heterogeneous: 16 clients having 3 kinds of NLU datasets (4 with SST-2, 7 with QNLI, 5 with QQP); and (3) Most heterogeneous: 16 clients having 4 kinds of NLU datasets (3 with SST2, 6 with QNLI, 2 with MRPC, 5 with QQP). The results in Table˜8 demonstrate that FedLEASE consistently outperforms baseline methods across all heterogeneity levels, with the performance advantage becoming more pronounced as heterogeneity increases.
@@
We further conduct the additional experiments under both task and label non-i.i.d. setting, and the label distribution is illustrated in Figure˜9. Note the task distribution is same as what we used in the main experiment. As shown in Table˜9, our proposed mthod still outperforms other baselines in this both task and label Non-IID setting.
Table 9: Performance under Task and Label Non-IID.
@@
Then we perform additional experiments under label non-i.i.d. setting and follow the exact setup used in FedSA under a Dirichlet ($\alpha=0.5$) partitionaing scheme using 20 clients on the QQP dataset. Results in Tables˜11, 13, 13 and 14 demonstrate that FedLEASE consistently outperforms existing methods even under label-heterogeneous conditions, further confirming the robustness and generality of our proposed approach.
Table 11: Label Distribution under Dirichlet(0.5) on QQP
Client ID 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
Class 0 (%) 0.7 18.1 70.3 8.0 0.4 85.8 92.3 18.7 90.9 33.2 99.9 20.6 81.8 85.9 24.6 9.8 80.5 92.1 93.1 75.2
Class 1 (%) 99.3 81.9 29.7 92.0 99.6 14.2 7.7 81.3 9.1 66.8 0.1 79.4 18.2 14.1 75.4 90.2 19.5 7.9 6.9 24.8
Table 12: Silhouette Scores under Only Label Non-IID
Clusters 2 3 4 5 6 7 8
Score 0.1829 0.1139 0.0553 0.0532 0.0482 0.0527 0.0506
Table 13: Cluster Groups under Only Label Non-IID
Group Clients
Group 0 2, 5, 6, 8, 10, 12, 13, 16, 17, 18, 19
Group 1 0, 1, 3, 4, 7, 9, 11, 14, 15
Table 14: Performance under Only Label Non-IID
Method FedIT FFA-LoRA FedDPA FedSA IFCA+LoRA FedLEASE
Accuracy (%) 83.52 83.05 85.78 86.85 84.73 89.23
Impact of Expert Upper Bound $M_{\max}$. In our main experiments, we set the expert upper bound to $M_{\max}=8$. To investigate the sensitivity of our approach to this parameter, we conducted additional experiments with lower upper bounds ($M_{\max}=2,3,4$) using the same experimental setup: 16 clients with data from 4 GLUE datasets. As shown in Table˜10, we observe comparable performance between $M_{\max}=8$ and $M_{\max}=4$ configurations. This aligns with our clustering analysis in Appendix˜B, which indicates that this particular configuration requires only 4 experts. This finding validates that our method can efficiently determine the appropriate number of experts needed even when the budget ($M_{\max}$) exceeds the system’s actual requirements. When restricted to $M_{\max}<4$, we observe a performance degradation compared to the $M_{\max}=4$ or $M_{\max}=8$ settings, further confirming the importance of allocating an adequate number of experts. Nevertheless, it is noteworthy that even with this constrained expert budget, our proposed method still outperforms all baseline methods. This demonstrates the robustness of our approach and its ability to make efficient use of even limited expert resources through effective allocation and adaptive selection.
Appendix D Computational Overhead
Our clustering step is performed only once during the initialization phase and is not repeated during iterative training. Thus, its runtime impact is negligible. As observed in Section˜3.3, using only the LoRA B matrices offers an efficient and lightweight proxy for task similarity, given their small size compared to full model weights or BA products. As shown in Table˜15, we measured the clustering time (3.11 seconds on Intel Xeon Platinum 8570 CPU), which is significantly shorter than the total training time (193.49 seconds with local training on the NVIDIA B200 GPU):
Table 15: Comparison of Computing Time
Time (s) FedIT FFA-LoRA FedDPA FedSA IFCA+LoRA FedLEASE
Local Per-Epoch Training Time 3.75 3.41 3.82 3.72 3.85 3.78
Global Aggregation Time 0.048 0.038 0.051 0.042 0.061 0.055
Clustering Time - - - - 2.77 3.11
Total Training Time 188.70 171.45 192.28 187.05 263.28 193.49
Appendix E Limitations
@@
Appendix F Convergence Analysis
@@ F.1 Assumptions
Assumption F.1 (Client-Level Smoothness).
@@
Assumption F.2 (Client-Level Strong Convexity).
@@
Assumption F.3 (Bounded Expert Sensitivity).
@@
Assumption F.4 (Cluster Assignment Stability).
@@
Theorem F.5 (Convergence of FedLEASE).